OpenStudy (anonymous):

Graph the following linear system to determine the number of solutions. y= -x+2 x+y=5

OpenStudy (anonymous):

First you want to put them both in slope intercept form. You will be able to look at the slopes. This is how you determine the number of solutions. The slope is the "m" in y=mx+b format. y=-x+2 has a slope of -1 y=-x+5 has the same slope of -1 This means they are both parallel. And the b values (y-intercepts) are different. This means they will be parallel lines next to each other and will never cross paths. Since intersections on graphs represent your solutions, the lack of an intersection means there will be no solution to this problem.

OpenStudy (anonymous):

for 1st eqn (2,0) & (0,2) 2nd (5,0) & (0,5) if we will plot it on graph it will be a parallel line !